# Emergence of a second law of thermodynamics in isolated quantum systems

### v1

@@ Appendix A Additional entropies
As we discussed in the main text, there is no unique entropy with respect to which one can examine irreversibility in the context of isolated quantum system’s evolution, and given different measures of entropy than the ones we used, one would have to adapt the results. For the main part of this work, we have focused on the Shannon entropy of an observable. The von Neumann entropy, however, is arguably the most commonly used entropy measure for quantum states, and thus, we discuss in Sec. I.1 how to understand the von Neumann entropy in the context of equilibration. In the context of observables the observational entropy is another common notion of entropy, and in Sec. I.2 we provide the results from our main text adapted to the observational entropy.
I.1 A second law for the von Neumann entropy?
@@
$\left|S_{\mathrm{vN}}\left[\langle\rho(t)\rangle_{T}\right]-S_{\mathrm{vN}}\left[\omega\right]\right|\leq\log(d)\vartheta+H_{2}[\vartheta],\text{ with }\vartheta=\frac{2\sqrt{d}}{\omega_{\mathrm{min}}T}$ (16)
and $\omega_{\mathrm{min}}:=\min\{|\lambda_{i}-\lambda_{j}|\neq 0\}$ is the smallest non-zero magnitude gap of $H$.
@@
$\left|S_{\mathrm{vN}}\left[\langle\rho(t)\rangle_{T}\right]-S_{\mathrm{vN}}\left[\omega\right]\right|\leq\frac{1}{2}\log(d)\|\langle\rho(t)\rangle_{T}-\omega\|_{1}+H_{2}\left[\frac{1}{2}\|\langle\rho(t)\rangle_{T}-\omega\|_{1}\right].$ (17)
@@
$=\sum_{ij}\left|\rho_{ij}\right|^{2}\left|\mathrm{sinc}\left(\frac{\omega_{ij}T}{2}\right)-\delta_{\omega_{ij}0}\right|^{2}$ (25)
$=\sum_{ij\,:\,\omega_{ij}\neq 0}\left|\rho_{ij}\right|^{2}\left|\mathrm{sinc}\left(\frac{\omega_{ij}T}{2}\right)\right|^{2}$ (26)
@@
which we can directly insert into the entropy inequality (17), to find the upper bound with $\vartheta=\frac{2\sqrt{d}}{\omega_{\mathrm{min}}T}$, as stated in Prop. I.1. Note that the statement in the given form only holds for values $\vartheta\leq 1/2$, again because $H_{2}$ is monotonously growing only for arguments $\leq 1/2$. To remedy this, one could replace $H_{2}$ by $\log 2$ for arguments $\vartheta>1/2$, to arrive at a more general statement. ∎
@@
I.2 Generalization for the observational entropy
@@
where again $p_{i}=\textnormal{Tr}[\Pi_{i}^{O}\rho]$ is the population of $\rho$ in the $i$th eigenspace of $O$ and $V_{i}=\textnormal{Tr}[\Pi_{i}^{O}]$ is the dimension of that eigenspace. The dimension $V_{i}$ can also be understood as the number of microstates that make up the $i$th macrostate of the observable $O$. By splitting up the sum into two parts using $\log\frac{p_{i}}{V_{i}}=\log p_{i}-\log V_{i},$ we can understand the observational entropy as a sum of the Shannon observable entropy and an averaged Boltzmann entropy,
@@
The first term, the Shannon observable entropy is what we have already discussed in the main text, corresponding to the entropy of the probabilistic distribution of the measurement outcomes. The second term, the averaged Boltzmann entropy, physically corresponds to an average of the Boltzmann entropy “$S=k_{B}\log W$” over the $r$ possible macroscopic outcomes of the observable with $V_{i}$ the number of microscopic states compatible with the $i$th outcome of the measurement of $O$. In Fig. 2 we provide a numerical example displaying how the observational entropy splits into the Shannon and Boltzmann terms. With the additional Boltzmann term, the observational entropy further coarse grains the information available in the state $\rho(t)$, and one may therefore ask whether this additional step is necessary to recover the second law relative to an observable.
@@
We answer this question negatively, pointing out that in some pathological cases the observational entropy even wrongly suggest that a system equilibrates with respect to an observable, when in fact, it does not. One example where this happens is in in subsystem equilibration. For illustration, we may consider a joint spin-$\frac{1}{2}$ and bath system where the bath is assumed to have dimension $d_{B}$. Suppose the spin is governed by the Hamiltonian $H_{S}=g\sigma_{x}$, it initially starts in the state $\left|\right.\!{0}\!\left.\right\rangle$, and the observable of interest is the spin polarization in $z$-direction, $O=\sigma_{z}\otimes\mathds{1}_{B}$. If the bath and the spin do not interact, the population vector for the measurement outcome evolves as $\vec{p}(t)=(\cos(gt)^{2},\sin(gt)^{2}),$ and therefore $S_{\mathrm{Sh}}^{O}[\rho(t)]$ periodically fluctuates between 00 and $\log 2$. The Boltzmann entropy term, on the other hand, is constant, $S_{B}^{O}[\rho(t)]=\log d_{B}$, and in the limit of a large bath, the relative fluctuations of the Boltzmann entropy vanish as $1/\log d_{B}$. Looking solely at the observational entropy, this suggests that the system equilibrates with respect to $O$, whereas physically the contrary is the case: the spin-$\frac{1}{2}$ subsystem under observation is isolated from the environment and does not equilibrate, as correctly captured by the Shannon observable entropy.
@@
In summary, this shows that the additional coarse graining of the observational entropy can result in cases where the Shannon entropy does not equilibrate while the observational entropy does, because of the inclusion of the Boltzmann term. Cases such as the one discussed before, however, are not physically self-consistent, because if one ignores fluctuations of order $\log 2$ from the observable, the operational way to model this is to use another observable that does not distinguish between those states, for example using a POVM instead of a PVM to model the fact that one can not differentiate between the spin orientations perfectly (see generalization in Appendix II.3). Operationally meaningful observable equilibration therefore requires the Shannon observable entropy to equilibrate, not the observational entropy. In the following we show that the condition $d_{\mathrm{eff}}\gg r$ under which the Shannon observable entropy equilibrates is also sufficient for the observational entropy to equilibrate in a similar sense:
@@
Proof..
We provide the proof of this statement in Sec. B together with the other proofs. ∎
@@
The origin for the $\log d$ scaling comes from the Boltzmann term which scales at worst with $\log d$, the maximum number of microstates in the system. However, due to the additional factor of $\eta_{\varepsilon,T}\propto\sqrt{{r}/{d_{\mathrm{eff}}}}$, if the effective dimension is of the order of the dimension of the system $d_{\mathrm{eff}}\sim d$, this additional correction will also vanish in the limit of large systems $d_{\mathrm{eff}}\gg r$. In a very close analogy to Cor. III, the observational entropy also satisfies a fluctuation theorem, bounding the probability of small fluctuations of $S_{\mathrm{Obs}}^{O}[\rho(t)]$ away from the equilibrium value $S_{\mathrm{Obs}}^{O}[\omega]$ as we show below:
@@
[Observable entropy fluctuations] Let $\rho(t),\rho^{\infty}$ and $\hat{O}$ be as in Lemma II.1. If $t\in\mathbb{R}_{\geq 0}$ is sampled uniformly randomly, we have
@@
$\nu=\log(d)\frac{r}{2\sqrt{d_{\mathrm{eff}}}}+g\left[\frac{r}{2\sqrt{d_{\mathrm{eff}}}}\right],$ (33)
with $g$ the function as in Prop. 2.
@@
The proof goes exactly as that of Cor. 2 with the difference that we define $X=\left|S_{\mathrm{Obs}}^{O}[\rho(t)]-S_{\mathrm{Obs}}^{O}[\omega]\right|$ for Chebyshev’s inequality. The reader is referred to the proof of Cor. III in Sec. B for details. ∎
Appendix B Proofs for entropy bounds
In this section, we prove all our main results and provide additional background to the statements and the formalism we use. In Sec. II.1 we provide the proofs for the average entropy difference bounds, that is Theorem 1 and Proposition 2. Following this, Sec. II.2 contains the proofs for the fluctuation bounds of Cor. III and Cor. 2.
II.1 Proofs for averaged entropy differences
Before discussion of the proof of Thm. 1, more details regarding the inequality (9) are supplied, in particular regarding the applicability of Jensen’s inequality: {lemma}[] Given $\rho(t)$, $\omega$ and $O$ as defined in the main text, we have in the infinite time-average,
Proof.
The proof of this statement is an application of the Jensen inequality which states that for a concave, real-valued function $g[X]$ the following inequality holds: $\left\langle g[X]\right\rangle\leq g\left[\left\langle X\right\rangle\right]$. In this context, $X$ is a real-valued random variable that can in general be defined on a convex subset of $\mathbb{R}^{r}$, as shown for example in Thm. 7.11 of [63]. For our setting, we have that $g[\,\cdot\,]=S_{\mathrm{Sh}}^{O}[\,\cdot\,]$ is concave as a function of the probability vector $\vec{p}(t)$ defined by the observable outcomes. Note that it is not concave as a function of $t$, though this is, of course, not needed here. Formally, the Shannon observable entropy as defined in the main text is a function of density matrices, but since it only depends on the population vectors, we simplify the notation here to simply write $S_{\mathrm{Sh}}^{O}[\vec{p}]$ to actually mean $S_{\mathrm{Sh}}^{O}[\rho]$, where $\vec{p}$ is the population vector of $\rho$ relative to the observable $O$. By identifying $X=\vec{p}(t)$, and taking the uniform measure over the space $t\in[0,T]$, Jensen’s inequality reads
[…]
Moving on towards proving our main result, Thm. 1, we devise the following proof strategy:
We start with the first part. Let us recall that $\vec{p}(t)=(p_{1},\dots,p_{r})$ with $p_{i}=\textnormal{Tr}[\Pi_{i}^{O}\rho]$ is the vector of populations of $\rho(t)$ with respect to the observable’s eigenspaces. Similarly, $\vec{p}_{\omega}$ is defined for the infinite time averaged state $\omega$. Then, in Eq. (10) of the main text we claimed that $\langle\|\vec{p}(t)-\vec{p}_{\omega}\|_{1}\rangle_{T}$ is small, which we prove in detail as part of Lemma II.1. The key reason why the populations’ differences are small on average is based on the results for observables from the theory of equilibration on average, which state that $\textnormal{Tr}[O\rho(t)]$ is close to $\textnormal{Tr}[O\omega]$ most of the time. One subtle point here concerns the number of macrostates: while for the observables’ expectation values it was sufficient to have that $d_{\mathrm{eff}}$ is large for equilibration, now, the the effective dimension has to be large compared to $r$, the number of macrostates corresponding to the observable $O$. We can therefore state that observable expectation values equilibrate under weaker conditions than the population vectors.
{lemma}
Proof.
$\left\langle\left|\textnormal{Tr}\left[\Pi_{i}^{O}\rho(t)\right]-\textnormal{Tr}\left[\Pi_{i}^{O}\omega\right]\right|^{2}\right\rangle_{T}\leq f(\varepsilon,T)\sqrt{\textnormal{Tr}\left[\Pi_{i}^{O\dagger}\Pi_{i}^{O}\omega^{2}\right]\textnormal{Tr}\left[\Pi_{i}^{O}\Pi_{i}^{O\dagger}\omega^{2}\right]}.$ (39)
[…]
$=f(\varepsilon,T)\textnormal{Tr}\left[\omega^{2}\right]$ (43)
$=\frac{f(\varepsilon,T)}{d_{\mathrm{eff}}}.$ (44)
Now that we have checked the first item of our proof strategy outlined in the beginning of Sec. B, we are ready to show Theorem 1, which we restate in the following for completeness.
See 1
Proof.
We subdivide the proof of this statement into two steps: (1) We express the left-hand side of Eq. (11) in terms of population differences, giving us a bound on the entropy difference on average expressed as a function of the average difference in populations between $\rho(t)$ and the infinite-time averaged state $\omega$. (2) Then, we use the result from Lemma II.1, which we can then insert into the bound found in the first step, finalizing the proof of this Proposition.
Step (1).
$\|\vec{\vartheta}\|_{1}=\frac{1}{2}\sum_{i=1}^{r}|\vartheta_{i}|,$ (46)
$\left\langle\left|S_{\mathrm{Sh}}^{O}[\rho(t)]-S_{\mathrm{Sh}}^{O}[\omega]\right|\right\rangle_{T}$ $=\left\langle\left|S_{\mathrm{Sh}}[\vec{p}(t)]-S_{\mathrm{Sh}}[\vec{p}_{\omega}]\right|\right\rangle_{T}$ (47)
$\leq\log(r-1)\langle\|\vec{\vartheta}\|_{1}\rangle_{T}+\langle H_{2}[\|\vec{\vartheta}\|_{1}]\rangle_{T}.$ (48)
$\left\langle\left|S_{\mathrm{Sh}}^{O}[\rho(t)]-S_{\mathrm{Sh}}^{O}[\omega]\right|\right\rangle_{T}\leq\log(r-1)\langle\|\vec{\vartheta}\|_{1}\rangle_{T}+H_{2}[\langle\|\vec{\vartheta}\|_{1}\rangle_{T}].$ (50)
Step (2).
Recalling that the binary entropy $H_{2}[x]$ is monotonously growing for $x\in[0,1/2]$ allows us to insert the inequality from Lemma II.1, $\langle\|\vec{\vartheta}\|_{1}\rangle_{T}\leq\eta_{\varepsilon,T}$ into the result from the first step (50). We find,
We now continue towards the bound for the observational entropy as introduced in Sec. I.2. The strategy for proving this result is related to that of Theorem 1, again relying on Lemma II.1, however, there are a few subtleties that come from the differences between the Shannon observable entropy and the observational entropy.
See 2
Proof.
Considering the observational entropy continuity bound derived in Theorem 6 of Ref. [64], with respect to the observable $O$, of two states $\rho$ and $\sigma$
$\big{|}S_{\mathrm{Obs}}^{O}[\rho]-S_{\mathrm{Obs}}^{O}[\sigma]\big{|}\leq g\left(\frac{1}{2}\|\rho-\sigma\|_{1}\right)+\frac{1}{2}\|\rho-\sigma\|_{1}\log(d),$ (52)
where $g(x)=-x\log(x)+(1+x)\log(1+x)$ and $d$ is the dimension of the Hilbert space $\mathcal{H}$ of $\rho$ and $\sigma$. On the other hand, it is important to notice that the observational entropy of a given $\rho$ is equal to the observational entropy of the coarse-grained state $\rho^{\mathrm{cg}}=\sum_{i=1}^{r}\textnormal{Tr}\left[\Pi_{i}^{O}\rho\right]\frac{\Pi_{i}^{O}}{V_{i}}$, with respect to a given observable $O=\sum_{i=1}^{r}O_{i}\Pi_{i}^{O}$,
$S_{\mathrm{Obs}}^{O}[\rho]=S_{\mathrm{Obs}}^{O}[\rho^{\mathrm{cg}}].$ (53)
Considering $\rho=\rho(t)$ and $\sigma=\omega$, it is possible to define two probability vectors with elements $\tilde{p}_{i}(t)=\textnormal{Tr}\left[\Pi_{i}\rho(t)\right]/V_{i}$ and $\tilde{p}_{\omega,i}=\textnormal{Tr}\left[\Pi_{i}\omega\right]/V_{i}$. In this case, the trace distance between the coarse-grained states is
$\|\rho(t)^{\mathrm{cg}}-\omega^{\mathrm{cg}}\|_{1}=\sum_{i}\|\tilde{p}_{i}(t)-\tilde{p}_{\omega,i}\|_{1}|V_{i}|=\sum_{i}|\textnormal{Tr}\left[\Pi_{i}\rho(t)\right]-\textnormal{Tr}\left[\Pi_{i}\omega\right]|=\sum_{i}|p_{i}(t)-p_{\omega,i}|.$ (54)
By Lemma II.1, the time average of the trace distance between the coarse-grain states for a given $\epsilon\geq 0$ and $T\geq 0$ is bounded as
$\Big{\langle}\|\rho(t)^{\mathrm{cg}}-\omega^{\mathrm{cg}}\|_{1}\Big{\rangle}\leq\eta_{\varepsilon,T}.$ (55)
From the time average of the continuity bound on Eq. (52), and by considering Eq. (53) and (55), we obtain
$\Big{\langle}\big{|}S_{\mathrm{Obs}}^{O}[\rho(t)]-S_{\mathrm{Obs}}^{O}[\omega]\big{|}\Big{\rangle}_{T}$ $=\Big{\langle}\big{|}S_{\mathrm{Obs}}^{O}[\rho(t)^{\mathrm{cg}}]-S_{\mathrm{Obs}}^{O}[\omega^{\mathrm{cg}}]\big{|}\Big{\rangle}_{T}\leq\log(d)\eta_{\varepsilon,T}+\Big{\langle}g\left(\frac{\|\rho(t)^{\mathrm{cg}}-\omega^{\mathrm{cg}}\|_{1}}{2}\right)\Big{\rangle}_{T},$ (56)
$\leq\log(d)\eta_{\varepsilon,T}+g\left(\frac{\Big{\langle}\|\rho(t)^{\mathrm{cg}}-\omega^{\mathrm{cg}}\|_{1}\Big{\rangle}_{T}}{2}\right)$ (57)
$\leq\log(d)\eta_{\varepsilon,T}+g\left(\eta_{\varepsilon,T}\right),$ (58)
where in the second and third lines, we use the concavity and monotonicity of $g(x)=-x\log x+(1+x)\log(1+x)$, proving the statement. ∎
II.2 Proofs for fluctuation theorems
Here we provide a proof for the fluctuation bounds for the Shannon observable entropy (Cor. III), which can also by straightforward analogy serve as the proof of Cor. 2 concerning the fluctuations of the observational entropy.
See III
Proof.
We provide additional details for the proof of Cor. III by showing a simple proof for the generalized Chebyshev inequality, which we recall here for completeness:
$P[X\geq\kappa]\leq\frac{E[g(X)]}{g(\kappa)}.$ (59)
[…]
$E[X]=\log(1+r)\frac{r}{2\sqrt{d_{\mathrm{eff}}}}+H\left[\frac{r}{2\sqrt{d_{\mathrm{eff}}}}\right]=:\delta,$ (63)
as a direct consequence of Prop. 1. Setting $\kappa:=\sqrt{\delta}$ yields
which completes the proof as claimed in the main text. ∎
II.3 Replacing projective observables with POVMs

### v2

@@ Appendix B Additional entropies
As we discussed in the main text, there is no unique entropy with respect to which one can examine irreversibility in the context of isolated quantum system’s evolution, and given different measures of entropy than the ones we used, one would have to adapt the results. For the main part of this work, we have focused on the Shannon entropy of an observable. The von Neumann entropy, however, is arguably the most commonly used entropy measure for quantum states, and thus, we discuss in Sec. II.1 how to understand the von Neumann entropy in the context of equilibration. In the context of observables the observational entropy is another common notion of entropy, and in Sec. II.2 we provide the results from our main text adapted to the observational entropy.
II.1 A second law for the von Neumann entropy?
@@
$\left|S_{\mathrm{vN}}\left[\langle\rho(t)\rangle_{T}\right]-S_{\mathrm{vN}}\left[\omega\right]\right|\leq\log(d)\vartheta+H_{2}[\vartheta],$ (42)
with $\vartheta={2\sqrt{d}}/{(\omega_{\mathrm{min}}T)}$ and $\omega_{\mathrm{min}}:=\min\{|\lambda_{i}-\lambda_{j}|\neq 0\}$ is the smallest non-zero magnitude gap of $H$.
@@
$\left|S_{\mathrm{vN}}\left[\langle\rho(t)\rangle_{T}\right]-S_{\mathrm{vN}}\left[\omega\right]\right|$ (43)
$\quad\leq\frac{1}{2}\log(d)\|\langle\rho(t)\rangle_{T}-\omega\|_{1}+H_{2}\left[\frac{1}{2}\|\langle\rho(t)\rangle_{T}-\omega\|_{1}\right].$
@@
$=\sum_{ij}\left|\rho_{ij}\right|^{2}\left|\mathrm{sinc}\left(\frac{\omega_{ij}T}{2}\right)-\delta_{\omega_{ij}0}\right|^{2}$
$=\sum_{ij\,:\,\omega_{ij}\neq 0}\left|\rho_{ij}\right|^{2}\left|\mathrm{sinc}\left(\frac{\omega_{ij}T}{2}\right)\right|^{2}$
@@
which we can directly insert into the entropy inequality (43), to find the upper bound with $\vartheta=\frac{2\sqrt{d}}{\omega_{\mathrm{min}}T}$, as stated in Proposition II.1. Note that the statement in the given form only holds for values $\vartheta\leq 1/2$, again because $H_{2}$ is monotonously growing only for arguments $\leq 1/2$. To remedy this, one could replace $H_{2}$ by $\log 2$ for arguments $\vartheta>1/2$, to arrive at a more general statement. ∎
@@
II.2 Generalization for the observational entropy
@@
where again $p_{i}=\textnormal{Tr}[\Pi_{i}^{O}\rho]$ is the population of $\rho$ in the $i$th eigenspace of $O$ and $V_{i}=\textnormal{Tr}[\Pi_{i}^{O}]$ is the dimension of that eigenspace. The dimension $V_{i}$ can also be understood as the number of microstates that make up the $i$th macrostate of the observable $O$. The observational entropy can also be understood as the entropy of the coarse-grained state $\rho^{\rm cg}=\sum_{i=1}^{r}p_{i}{\Pi_{i}^{O}}/{V_{i}}$ (note that this does not directly generalize to POVMs, as opposed to the remaining results) [8, 52],
$S_{\rm vN}[\rho^{\rm cg}]$ $=-\sum_{i=1}^{r}\textnormal{Tr}\left[p_{i}\frac{\Pi_{i}^{O}}{V_{i}}\log\frac{p_{i}}{V_{i}}\right]$ (53)
$=S_{\rm Obs}^{O}[\rho].$ (54)
Contrary to the Shannon observable entropy that captures the knowledge attainable purely from the measurement outcome, we can therefore understand the observational entropy as the entropy associated with the system, given access knowledge of only the coarse-grained state.
By splitting up the sum into two parts using $\log\frac{p_{i}}{V_{i}}=\log p_{i}-\log V_{i},$ we can understand the observational entropy as a sum of the Shannon observable entropy and an averaged Boltzmann entropy,
@@
The first term, the Shannon observable entropy is what we have already discussed in the main text, corresponding to the entropy of the probabilistic distribution of the measurement outcomes. The second term, the averaged Boltzmann entropy, physically corresponds to an average of the Boltzmann entropy “$S=k_{B}\log W$” over the $r$ possible macroscopic outcomes of the observable with $V_{i}$ the number of microscopic states compatible with the $i$th outcome of the measurement of $O$. In Fig. 3 we provide a numerical example displaying how the observational entropy splits into the Shannon and Boltzmann terms. With the additional Boltzmann term, the observational entropy further coarse grains the information available in the state $\rho(t)$, and one may therefore ask whether this additional step is fundamentally necessary to recover the second law relative to an observable.
@@
We answer this question negatively, by first of all noting that the ratio $r/d_{\rm eff}\ll 1$ is sufficient for equilibration of both the Shannon observable entropy as well as the observational entropy (Proposition 3). Secondly, there are some pathological cases where the observational entropy even wrongly suggest that a system equilibrates with respect to an observable, when in fact, it does not. One example where this happens is in in subsystem equilibration. For illustration, we may consider a joint spin-$\frac{1}{2}$ and bath system where the bath is assumed to have dimension $d_{B}$. Suppose the spin is governed by the Hamiltonian $H_{S}=g\sigma_{x}$, it initially starts in the state $\left|\right.\!{0}\!\left.\right\rangle$, and the observable of interest is the spin polarization in $z$-direction, $O=\sigma_{z}\otimes\mathds{1}_{B}$. If the bath and the spin do not interact, the population vector for the measurement outcome evolves as $\vec{p}(t)=(\cos(gt)^{2},\sin(gt)^{2}),$ and therefore $S_{\mathrm{Sh}}^{O}[\rho(t)]$ periodically fluctuates between 00 and $\log 2$. The Boltzmann entropy term, on the other hand, is constant, $S_{B}^{O}[\rho(t)]=\log d_{B}$, and in the limit of a large bath, the relative fluctuations of the Boltzmann entropy vanish as $1/\log d_{B}$. Looking solely at the observational entropy, this suggests that the system equilibrates with respect to $O$, whereas physically the contrary is the case: the spin-$\frac{1}{2}$ subsystem under observation is isolated from the environment and does not equilibrate, as correctly captured by the Shannon observable entropy.
@@
In summary, this shows that the additional coarse graining of the observational entropy can result in cases where the Shannon entropy does not equilibrate while the observational entropy does, because of the inclusion of the Boltzmann term. Cases such as the one discussed before, however, are not physically self-consistent, because if one ignores fluctuations of order $\log 2$ from the observable, the operational way to model this is to use another observable that does not distinguish between those states, for example using a POVM instead of a PVM to model the fact that one can not differentiate between the spin orientations perfectly (see generalization in SM Ref. C). Observable equilibration in an operational sense therefore requires the Shannon observable entropy to equilibrate, not necessarily the observational entropy. The reason for this apparent difference is that the observational entropy captures the entropy of the system given the information accessible via the observable (as per eq. (54)), whereas the Shannon observable entropy captures the information of the observable. In the following we show that the condition $d_{\mathrm{eff}}\gg r$ under which the Shannon observable entropy equilibrates is also sufficient for the observational entropy to equilibrate in a similar sense:
@@
To prove the desired inequality for the observational entropy, we want to proceed similarly to the proof of Theorem 2. In contrast to the Shannon entropy, however, we can not directly use inequality (27) to relate the difference in observational entropies to differences in the populations relative to the observable $O$. The following Lemma, which is a tightening of the continuity bound in Theorem 6 from Ref. [66], provides the desired statement. {lemma}[Observational entropy continuity bound] Let $\rho$ and $\sigma$ be two quantum states defined on a $d$-dimensional Hilbert space, and have $O$ be an observable on the same space. Define $\vec{p}_{\rho}$ and $\vec{p}_{\sigma}$ to be the population vectors of $\rho$ and $\sigma$, respectively, relative to $O$. Then,
$|S_{\mathrm{Obs}}^{O}[\rho]-S_{\mathrm{Obs}}^{O}[\sigma]|\leq g(\|\vec{p}_{\rho}-\vec{p}_{\sigma}\|_{1})+\log d\,\|\vec{p}_{\rho}-\vec{p}_{\sigma}\|_{1},$ (57)
where $g(x)=-x\log(x)+(1+x)\log(1+x)$ as before, and the 1-norm $\|\cdot\|_{1}$ for vectors is defined as in the main text (or equivalently eq. (28) of the Appendix).
Proof.
Let us start by considering a function $Z[\rho]$ defined on finite dimensional quantum states $\rho$ which is bounded concave. Bounded convacity means that for any convex combination $\rho=\sum_{k}\lambda_{k}\rho_{k},$ with $\lambda_{k}\geq 0,$ and $\sum_{k}\lambda_{k}=1$ and $\rho_{k}$ a well-defined quantum state, we have [66]
$0\leq Z\left[\rho\right]-\sum_{k}\lambda_{k}Z[\rho_{k}]\leq S_{\rm Sh}[\{\lambda_{k}\}_{k}],$ (58)
where $S_{\rm Sh}[\{\lambda_{k}\}_{k}]=-\sum_{k}\lambda_{k}\log\lambda_{k}$ is the Shannon entropy of the distribution $\{\lambda_{k}\}_{k}$. It has been shown in [66], that any such function satisfies the inequality
$|Z[\rho]-Z[\sigma]|\leq\kappa\|\rho-\sigma\|_{1}+g(\|\rho-\sigma\|_{1}),$ (59)
where $\kappa:=\sup_{\mu,\nu}|Z[\mu]-Z[\nu]|$. Note the difference in normalization of the trace norm in our work eq. (44) to that of [66]. For our purposes let us consider an $r$-dimensional state-space where we identify the population vector $\vec{p}_{\rho}$ with a quantum state
$\vec{p}_{\rho}=\sum_{i=1}^{r}\textnormal{Tr}\left[\Pi_{i}^{O}\rho\right]|{i}\rangle\!\langle{i}|,$ (60)
and similarly for $\vec{p}_{\sigma}$. We now define $Z[\,\cdot\,]$ on this $r$-dimensional state space by setting
$Z[\tau]:=-\sum_{i=1}^{r}\tau_{ii}\log\frac{\tau_{ii}}{V_{i}},$ (61)
where $\tau_{ii}$ is the diagonal entry of the $r\times r$ quantum state $\tau$. Inserting the diagonal state $\vec{p}_{\rho},$ we recover the observational entropy $Z[\vec{p}_{\rho}]=S_{\rm Obs}^{O}[\rho]$. We note that $Z$ is bounded concave: let $\tau=\sum_{k}\lambda_{k}\tau_{k}$ be a finite convex combination, then
$Z[\tau]-\sum_{k}\lambda_{k}Z[\tau_{k}]$ $=S_{\rm Sh}[\{\tau_{ii}\}_{i}]+\sum_{i=1}^{r}\tau_{i}\log V_{i}-\sum_{k}\lambda_{k}S_{\rm Sh}[\{\tau_{k,ii}\}_{i}]-\sum_{k}\sum_{i=1}^{r}\lambda_{k}\tau_{k,ii}\log V_{i}$ (62)
$=S_{\rm Sh}[\{\tau_{ii}\}_{i}]-\sum_{k}\lambda_{k}S_{\rm Sh}[\{\tau_{k,ii}\}_{i}],$ (63)
is equal the difference of the diagonal Shannon entropies. The Shannon entropy is bounded concave by Lemma 1 of Ref. [66], and thus, also $Z$ is. Therefore, our definition (61) satisfies all the necessary assumptions for the inequality (59) to hold. By applying this inequality to the diagonal states $\vec{p}_{\sigma}$ and $\vec{p}_{\rho}$ we find that
$|Z[\vec{p}_{\rho}]-Z[\vec{p}_{\sigma}]|\leq\kappa\|\vec{p}_{\rho}-\vec{p}_{\sigma}\|_{1}+g(\|\vec{p}_{\rho}-\vec{p}_{\sigma}\|_{1}).$ (64)
Note that because $\vec{p}_{\rho}$ and $\vec{p}_{\sigma}$ are diagonal in the same basis, the 1-norm for vectors defined in eq. (28) and the 1-norm for states defined in eq. (44) agree. Using that $\kappa=\log d$ for the observational entropy, we conclude the proof.
Note that this result, heavily based on the work in Ref. [66], is a tightening of Theorem 6 from the same reference. The underlying reason is that the trace-distance between two quantum states $\rho$ and $\sigma$ is always lower bounded by the 1-norm distance of their populations relative to any observable [62],
$\|\rho-\sigma\|_{1}\geq\|\vec{p}_{\rho}-\vec{p}_{\omega}\|_{1}.$ (65)
Thus, the righthand side of (57) is always smaller that of Theorem 6 from [66]. With this bound at hand, we are ready to prove Proposition 3.
Proof of Proposition 3..
Applying Lemma 3 to this setting, we find that
$\Big{\langle}\big{|}S_{\mathrm{Obs}}^{O}[\rho(t)]-S_{\mathrm{Obs}}^{O}[\omega]\big{|}\Big{\rangle}_{T}$ (66)
$\quad\leq\Big{\langle}\log d\,\|\vec{p}(t)-\vec{p}_{\omega}\|_{1}+g\left(\|\vec{p}(t)-\vec{p}_{\omega}\|_{1}\right)\Big{\rangle}_{T}$ (67)
$\quad\leq\log d\,\eta_{\varepsilon,T}+g\left(\eta_{\varepsilon,T}\right),$ (68)
where in the second line, we use the concavity and monotonicity of $g(x)=-x\log x+(1+x)\log(1+x)$, proving the statement. ∎
@@
The origin for the $\log d$ scaling comes from the Boltzmann term which scales at worst with $\log d$, the maximum number of microstates in the system. However, due to the additional factor of $\eta_{\varepsilon,T}\propto\sqrt{{r}/{d_{\mathrm{eff}}}}$, if the effective dimension is of the order of the dimension of the system $d_{\mathrm{eff}}\sim d$, this additional correction will also vanish in the limit of large systems $d_{\mathrm{eff}}\gg r$. In a very close analogy to Corollary III, the observational entropy also satisfies a fluctuation theorem, bounding the probability of small fluctuations of $S_{\mathrm{Obs}}^{O}[\rho(t)]$ away from the equilibrium value $S_{\mathrm{Obs}}^{O}[\omega]$ as we show below:
@@
[Observable entropy fluctuations] Let $\rho(t),\rho^{\infty}$ and $\hat{O}$ be as in Lemma I.1. If $t\in\mathbb{R}_{\geq 0}$ is sampled uniformly randomly, we have
@@
$\nu=\frac{\log(d)}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}+g\left[\frac{1}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}\right],$ (70)
with $g$ the function as in Proposition 3.
@@
The proof goes exactly as that of Corollary 3 with the difference that we define $X=\left|S_{\mathrm{Obs}}^{O}[\rho(t)]-S_{\mathrm{Obs}}^{O}[\omega]\right|$ for Chebyshev’s inequality. The reader is referred to the proof of Corollary III in Sec. A for details. ∎
Akin to the generalization in eq. (40) of the fluctuation bounds for the Shannon observable entropy from Corollary III, we can also generalize Corollary 3 for the case of the observational entropy,
$P\left[\left|S_{\mathrm{Obs}}^{O}[\rho(t)]-S_{\mathrm{Obs}}^{O}[\omega]\right|\geq\mu\right]\leq\frac{\nu}{\mu}.$ (71)
As before, $\mu$ is the small parameter relative to which we compare the observational entropy fluctuations on the left-hand side.
Appendix C Replacing projective observables with POVMs